It is hot a day and Bert is thirsty ,here is the value he places on a bottle of water
Value of first 7$
Value of second bottle 5$
Value of third bottle 3$
Value of fourth bottle 1$
a)from this information derive Bert,s demand schedule graph his demand curve for bottled water.
b)if the price of bottle of water is $4 how many bottle does Bert buy?how much consumer surplus does Bert get from his purchases ? Show Bert,s consumer surplus in your graph.
c)if the price fall to$2 how does quantity demanded change how does Bert's consumer surplus change show these changes in your graph
1
Expert's answer
2020-06-24T16:37:13-0400

a)from this information derive Bert,s demand schedule graph his demand curve for bottled water.







b)if the price of bottle of water is $4 how many bottle does Bert buy?how much consumer surplus does Bert get from his purchases ? Show Bert,s consumer surplus in your graph.


If the price is $4, Bert will consume 2.5 bottles. The amount of the consumer surplus is:


"CS = \dfrac{1}{2}(9 - 4)(2.5) = \$6.25"




The triangular area is the consumer surplus.


c)if the price fall to$2 how does quantity demanded change how does Bert's consumer surplus change show these changes in your graph


If the price reduces to $2, the quantity demanded increases to 3.5 bottles.


The consumer surplus increases to:


"CS = \dfrac{1}{2}(9 - 2)(3.5) = \$12.25"